Use the given data to find the minimum sample size required to estimate the population proportion.

Margin of error: 0.09; confidence level: 90%; from a prior study,  (p-hat) is estimated by 0.17.
Answer #1

z value for 90% CI is 1.645 as P(-1.645<z<1.645)=0.90

So we will find n using formula of E

E=z*\sqrt{\frac{pq}{n}}

Hence n=(\frac{z}{E})^2*p*q=(\frac{1.645}{0.09})^2*0.17*0.83=47.17=48